### v1

# MiraGe: Editable 2D Images using Gaussian Splatting

Abstract
Implicit Neural Representations (INRs) approximate discrete data through continuous functions and are commonly used for encoding 2D images. Traditional image-based INRs employ neural networks to map pixel coordinates to RGB values, capturing shapes, colors, and textures within the network’s weights. Recently, GaussianImage has been proposed as an alternative, using Gaussian functions instead of neural networks to achieve comparable quality and compression. Such a solution obtains a quality and compression ratio similar to classical INR models but does not allow image modification. In contrast, our work introduces a novel method, MiraGe, which uses mirror reflections to perceive 2D images in 3D space and employs flat-controlled Gaussians for precise 2D image editing. Our approach improves the rendering quality and allows realistic image modifications, including human-inspired perception of photos in the 3D world. Thanks to modeling images in 3D space, we obtain the illusion of 3D-based modification in 2D images. We also show that our Gaussian representation can be easily combined with a physics engine to produce physics-based modification of 2D images. Consequently, MiraGe allows for better quality than the standard approach and natural modification of 2D images.
1 Introduction
Figure 1: MiraGe encodes 2D images with parameterized Gaussians, enabling high-quality reconstruction and real-life-like modifications. Selected part of image can be transformed in 3D space, creating a 3D effect, with a physics engine controlling movement and interactions.
Recent research has increasingly emphasized human perception and the understanding of the world through this lens (Lu, 2019; Davoodi et al., 2023). In line with this trend, we introduce a model that encodes 2D images by simulating human interpretation. Specifically, our model perceives a 2D image as a human would view a photograph or a sheet of paper, treating it as a flat object within a 3D space. This approach allows for intuitive and flexible image editing, capturing the nuances of human perception while enabling complex transformations (see Fig. 1).
Gaussian Splatting (3DGS) framework models the structure of a 3D scene using Gaussian components (Kerbl et al., 2023). In the 2D domain, GaussianImage (Zhang et al., 2024) has shown promising results in image reconstruction by efficiently encoding images in the 2D space, with a strong focus on model efficiency and reduced training time. Regrettably, the GaussianImage lacks the capability for human-driven adjustments of scene objects, which is a crucial aspect of 3DGS. In our work, we address this by introducing the MiraGe [^1] https://github.com/waczjoan/MiraGe/blob/main/README.md model, which encodes 2D images through the lens of human perception, bridging the gap between 2D image representation and 3D spatial understanding (see Fig. 2).
Figure 2: MiraGe employs 3D flat parameterized Gaussians in 3D space to encode 2D images, representing each flat Gaussian as three points, forming a cloud of triangles called a triangle soup. This representation enables real-time manipulation of the 3D triangle/point clouds, allowing for flexible, real-world modifications. The model seamlessly integrates with a physics engine, enhancing its applicability in dynamic environments.
Building on the foundational idea that humans intuitively can perform transformations on photographs–primarily through affine transformations and bending them beyond the 2D plane–we introduce a novel approach using flat Gaussians with GaMeS parameterization (Waczyńska et al., 2024b). This allows our model to facilitate image editing within 2D and 3D spaces. In particular, our framework enables changes in the image’s perspective with minimal effort. Perspective adjustments, typically complex, are simplified by direct editing in the third dimension (see Fig. 3).
Figure 3: MiraGe simplifies intuitive editing of image, allowing transformations such as adjusting the tilt of a hand with minimal complexity. This is achieved by modifying the object along the third dimension.
In addition to classical edits, our model has the unique capability of interfacing with physics engines, enabling applications that enhance the realism and immersiveness of animations (Jiang et al., 2024). MiraGe treats the physics engine as a black box and offers three distinct methods for controlling Gaussians, i.e., 2D, Amorphous and Graphite. For 2D representation (2D-MiraGe) we used Taichi_elements, for 3D representation (Amorphous-MiraGe, Graphite-MiraGe) we use Blender. This flexibility makes our model highly applicable to various fields, such as computer graphics for populating spatial interfaces, where realistic, physics-factual 2D animations can be incorporated (Tadeja et al., 2023).
Embedding 2D images in 3D space allows for seamless integration of 2D and 3D objects, enabling the creation of dynamic backgrounds or interactive elements within animated scenes. This versatility extends to applications such as virtual reality, where 2D images can function as backdrops (Yin et al., 2024). This capability opens up new avenues for creative composition, offering a powerful toolset for users. The novelty of this work lies in its ability to enable easy, intuitive 3D transformations and integrations within a traditionally 2D framework, expanding the possibilities for both image editing and animation (see Fig. 4).
Since high-quality image reconstruction is critical in animation, we compared MiraGe with other models, showing our state-of-the-art performance in the image reconstruction task. Our results demonstrate that the model operates in real-time, though this comes with a trade-off in terms of model size (see Fig. 5).
Figure 4: Two images were encoded using the MiraGe model on distinct planes within a 3D space. This setup allows for seamless integration of the encoded images, resulting in a collage-like composition. Moreover, the model facilitates editing capabilities, as illustrated here with modifications to the background image (the rear plane).
It is worth highlighting that flat 3D Gaussians can be utilized for 2D images in four distinct scenarios, with modifications that emphasize how controlling the Gaussians during training affects the perspective of viewing each image (see Fig. 6).
The following constitutes a list of our key contributions:
  • We introduce the MiraGe model, which represents 2D images using flat Gaussian components, achieving state-of-the-art reconstruction quality.
  • MiraGe enables the manipulation of 2D images within 3D space, creating the illusion of 3D effects.
  • We integrate MiraGe with a physics engine, enabling physics-based modifications and interactions for both 2D and 3D environments
Figure 5: MiraGe outperforms other models in the image reconstruction task. The following columns present ground truth image, GaussianImage reconstruction and MiraGe reconstruction. The bottom row graphics show the difference between the ground truth image and the outcomes of subsequent methods.
2 Related Works
Our work builds on several key research areas, including image reconstruction techniques, Gaussian-based representations and Gaussian animation frameworks.
One rapidly growing area in image reconstruction is Implicit Neural Representations (INRs), which have attracted significant attention for their ability to model continuous signals, such as images, through neural networks (Klocek et al., 2019). INRs encode spatial coordinates and map them to corresponding values, such as RGB color, allowing for highly compact and efficient representations (Xie et al., 2022). This has led to the development of several specialized models for image INR, such as SIREN (Sitzmann et al., 2020a) with the novelty of sine used as a periodic activation function to tackle the problem of complex image signals. Fourier feature mapping was proposed in (Tancik et al., 2020) as another answer to the difficulty of aligning the multilayer perceptron (MLP) predictions with high-frequency pictures. Interestingly, authors of WIRE (Saragadam et al., 2023a) have leveraged continuous complex Gabor wavelets to capture visual signals with decent quality. The growing field of research resulted in further improvements of already existing solutions, e.g. in (Liu et al., 2024), certain limitations of SIREN, namely the arising capacity-convergence gap, were successfully alleviated with the idea of variable-periodic activation functions. Yet another worth noting work from this area is (Müller et al., 2022) with INR solution designed to effectively perform on modern computer architecture utilizing a simple data structure concept of hashmap to offer speed-oriented image representation with high fidelity of the outcomes.
Alternative approaches to INRs were presented in GaussianImage (Zhang et al., 2024). Instead of natural networks, the authors propose to approximate 2D images using 2D Gaussian components. In practice, GaussianImage is a 2D version of 3DGS (Kerbl et al., 2023) that uses 2D Gaussians instead of their 3D version and a simplified rendering procedure. Thanks to such a modification, the GaussianImage is invariant to the order of Gaussian components. Therefore, such a model is numerically efficient.
Figure 6: Flat 3D Gaussians can be utilized for 2D images in four distinct scenarios. Modifications demonstrate how the perspective of viewing a 2D image can change in each case. We establish a baseline using the GaMeS framework with Gaussian initialization on the XZ plane. Our results show that controlling Gaussian positions during learning enables various applications. The model employs a mirror setup during training, with the Amorphous configuration achieving the best results for image reconstruction and 3D analysis. The 2D model represents images on a single plane, allowing 2D physics engines like Taichi_elements to be used, but it does not support 3D modifications. In contrast, the Graphite model operates across multiple planes, making it ideal for 3D spatial reasoning and image combination.
GaussianImage represents each pixel color as a weighted sum of 2D Gaussians. The training procedure is similar to 3DGS without pruning. The authors show that such representation gives a similar reconstruction quality to classical INR models and is able to obtain a high compression ratio and fast rendering.
The representation and editing of objects using Gaussians is a well-explored topic in 3D graphics. In this field, meshes can be modified to simulate Gaussian editing (Guédon & Lepetit, 2024; Huang et al., 2024), or Gaussians can be directly parameterized and manipulated to achieve specific outcomes (Waczyńska et al., 2024b; a). This approach enables flexible and continuous deformations, offering an intuitive method for controlling object shapes and rendering properties, which has proven particularly useful in tasks like texture mapping, surface smoothing, and dynamic simulations.
Gaussians enable precise and flexible editing of objects, providing continuous control over shapes and transformations. Moreover, integrating physics engines enhances these capabilities, allowing for more sophisticated and physically consistent modifications, such as simulating realistic interactions, deformations and movements in 3D environments. (Xie et al., 2024; Borycki et al., 2024).
3 MiraGe: Editable 2D Images using Gaussian Splatting
Here, we describe in detail the inner workings of our MiraGe model. We start by presenting classical 3DGS. Next, we present GaMeS-based (Waczyńska et al., 2024b) parametrization of flat Gaussians. In the end, we present our MiraGe and how it relates to prior works.
Figure 7: MiraGe use GaMeS (Waczyńska et al., 2024b) representations of flat Gaussian by triangle soup. Therefore, we can use real-life modification by moving points.
3D Gaussian Splatting
3DGS models 3D scene by a set of Gaussian components with color and opacity:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},\Sigma_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$
defined by their mean (position) ${\bf m}_{i}$, covariance matrix $\Sigma_{i}$, opacity $\sigma_{i}$, and color $c_{i}$, which is represented using spherical harmonics (SH) (Fridovich-Keil et al., 2022).
During the rasterization stage, the 3DGS produces a sorted Gaussian list based on the projected depth information. Then, the $\alpha$-blending method is used to create the image. We refine the Gaussian parameters, color, and opacity in the training phase according to the reconstruction cost function. Training the Gaussian component is challenging. Consequently, the authors implement additional strategies for reducing and multiplying components. Gaussians with low opacity are removed, while those that change rapidly during optimization are multiplied. These strategies make the 3D Gaussian approach very efficient and capable of generating high-quality renders.
Figure 8: We integrate MiraGe with Material Point Methods (MPM) to achieve realistic alterations of 2D images. The initial column presents the original image, the subsequent two columns display renders captured midway through the simulation, and the final column shows the outcome of the full simulation. The colored lines in the last column trace the paths of randomly chosen points from the simulation.
GaMeS Parametrization of Gaussian Component
In MiraGe, we use flat Gaussian components in 3D space. In such a model we use Gaussian components with a covariance matrix $\Sigma$, factored as:
$\Sigma=RSSR^{T},$ (1)
where $R$ is the rotation matrix, and $S$ is a diagonal matrix containing the scaling parameters. However, we force one of the scale parameters to be zero. Consequently, we obtain a collection of flat Gaussian:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},R_{i},S_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (2)
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$. In such a case, we can use GaMeS (Waczyńska et al., 2024b) parametrization to represent flat Gaussian by triangle-face mesh. This mapping is denoted by $\mathcal{T}(\cdot)$. When applied, this parametrization generates a set of triangles labeled as triangle soup.
To outline the GaMeS parameterization, consider a Gaussian component $\mathcal{N}({\bf m},R,S)$, characterized by the mean ${\bf m}$, the rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and the scaling matrix $S=\mathrm{diag}(\varepsilon,s_{2},s_{3})$. Then its face representation $\mathcal{N}(V)$ is based on a triangle:
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ (3)
with the vertices defined as: ${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
$\mathcal{N}(\hat{\bf m},\hat{R},\hat{S})=\mathcal{N}(\mathcal{T}^{-1}(V))$ (4)
through the mean $\hat{\bf m}$, the rotation matrix $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$, and the scaling matrix $\hat{S}=\mathrm{diag}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3})$, where the parameters are defined by the following formulas:
$\hat{\bf m}={\bf v}_{1},\;\;\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$ (5)
$\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|},\;\;\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}),$ (6)
$s_{1}=\varepsilon,\;\;\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|,\;\;\textrm{and}\;\;\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle.$ (7)
Here $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process (Björck, 1994). Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as $\hat{\Sigma}=\hat{R}\hat{S}\hat{S}\hat{R}^{T}$.
Figure 9: MiraGe can be integrated with Blender, by using flat 3D Gaussians in 3D space. The initial column presents the original image, the subsequent two columns display renders captured midway through the simulation, and the final column shows the outcome at the simulation’s conclusion. The colored lines in the last column trace the paths of 10 randomly chosen points from the simulation.
The parametrization enables control over the Gaussians’ position, scale, and rotation by manipulating the underlying triangle mesh. Applying transformations to the triangle directly alters the corresponding Gaussian, as illustrated in Fig. 7.
MiraGe
Figure 10: MiraGe enables modifying 2D images, such as adjusting the scene’s elements’ sizes.
In this work, we present an approach that leverages the concept of flat Gaussian distributions in 3D space to model a single 2D image as input. Our methodology is grounded in human visual perception. This perspective allows us to reframe the problem: instead of merely processing a pixel matrix, we interpret the images as objects with a fixed spatial configuration in a 3D environment.
We put the 2D image on the $XZ$ plane where the center is situated at axes origin $(0,0,0)$ with the fixed distance from the camera origin. In practice, the distance from the plane is a hyper-parameter. In our approach, we model flat objects within 3D space, where the camera distance parameter effectively controls the perceived scale of the object. This relationship allows for intuitive adjustments of object size based on the desired visual effect. For instance, increasing the camera distance can naturally expand the apparent size of background elements like distant mountains (Fig. 4), making it easier to represent them as larger objects without additional modeling complexity. While this feature is beneficial, it is not strictly necessary for most applications.
Figure 11: Comparison of PSNR obtained on a butterfly image from DIV2K dataset by MiraGe in comparison with GaussianImage (Zhang et al., 2024) and Gaussian Splatting (Kerbl et al., 2023). The lines with markers represent how the PSNR was changing during training. Different colors represent models trained with different numbers of Gaussians during initialization. The dashed black and orange lines represent the best results obtained during training by GaussianImage and GS, respectively. Vertical lines represent iteration, where MiraGe obtained better results than GaussianImage, the time in min:sec format above each line is the training time until this iteration. The experiment was performed on the rtx 4070 GPU.
We propose a method that situates the Gaussians within the $XZ$ plane, ensuring that the entire image remains visible under perspective projection. To achieve this, the possible range of $x$-values and $z$-values is calculated using the camera field of view. We first calculate the deviation from 00 on the $X$ axis using the similarity of triangles $\text{dev}_{z}=\text{cam}_{\text{dist}}\cdot\tan(0.5\cdot\text{Fov}_{\text{vert}})$, where $\text{cam}_{\text{dist}}$ and $\text{Fov}_{\text{vert}}$ are camera distance from the $XZ$ plane and camera field of view respectively. The deviation in the $X$ axis can be then computed by multiplying this value by the camera aspect ratio.
Consequently, the initialization of Gaussians is consistently performed on the $XZ$ plane; however, we have opted to permit their movement within the 3D space. Drawing inspiration from three distinct models, we introduce three conceptual approaches for manipulating the spatial positioning of Gaussians.
Amorphous
The baseline approach how to control Gaussians is based on the classical GaMeS parametrization, initialized randomly on the $XZ$ plane, with the mean parameter’s $y$ coordinate set to zero:
$\mathcal{G}=\{(\mathcal{N}([m_{1},0,m_{3}],[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}],\mathrm{diag}(\varepsilon,s_{2},s_{3})),\sigma_{i},c_{i})\},$ (8)
where ${\bf m}=[m_{1},0,m_{3}]$ $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$.
It should be highlighted that we only initialized the Gaussian component on the $XZ$ plane. During training, Gaussians can move amorphously in 3D space. We use the classical loss function $L_{1}$ combined with a D-SSIM term:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}(I,GS(I))+\lambda\mathcal{L}_{D-SSIM}(I,GS(I)),$
where $I$ is the input image and $GS(I)$ is the constraint obtained by the Gaussian renderer. While this solution enables the modeling of images using a collection of triangles, often referred to as “triangle soup,” it proves insufficient for high-quality representations. During editing, significant artifacts emerge (Fig. 6–Baseline).
2D
Building on the promising outcomes of GaussianImage, we strategically anchored all Gaussians onto the $XZ$ plane. This configuration allows us to effectively translate the flat image geometry into a spatial framework, bridging perceptual intuition. We set the mean of these components to have zero in the second coordinate. Moreover, we use the projection of flat Gaussians on a 2D plane. Unfortunately, orthogonal projection can produce artifacts. Therefore, we use a rotation of Gaussian components to lay on the $XZ$ plane. Since we use flat Gaussians to extract such rotation we can use a rotation matrix between two vectors to align the vector in 3D (Markley, 1993). We use the notation $\mathrm{Rot}(a,b)$ for the rotation matrix.
MiraGe on 2D plane is define by set of 3G parameterized Gaussian components:
$\mathcal{G}=\{(\mathcal{N}([m_{1},0,m_{3}],[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),\mathrm{diag}(\varepsilon,s_{2},s_{3})),\sigma_{i},c_{i})\},$ (9)
where ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$.
Graphite
Figure 12: Comparison of FPS obtained on a butterfly image from DIV2K dataset by MiraGe in comparison with GaussianImage (Zhang et al., 2024) and Gaussian Splatting (Kerbl et al., 2023). The experiment was performed on the rtx 4070 GPU.
Unfortunately, 2D-MiraGe produces artifacts when we use modification in 3D space (see the third row in Fig. 6). Such an effect is coursed by the Gaussians, which appear randomly according to the camera position. To solve such a problem and obtain the possibility of 3D modifications, MiraGe allows the Gaussians to leave the $XZ$ plane:
$\mathcal{G}=\{(\mathcal{N}([m_{1},0,m_{3}]+\gamma{\bf e}_{2},[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),\mathrm{diag}(\varepsilon,s_{2},s_{3})),\sigma_{i},c_{i})\},$ (10)
where $\gamma$ is trainable parameter of translation scale along the vector ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as: $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$. Such a model allows for the order of Gassians according to camera positions.
By leveraging parameterized Gaussians, we achieved precise manipulation of 2D images directly within their native 2D space, enabling targeted edits of segmented regions and transformations of complete scenes in easier way. While this approach demonstrated substantial promise, we observed significant artifacts when extending manipulations into the 3D domain, particularly along the $Y$-axis.
To address this challenge, we employ a novel approach utilizing two opposing cameras positioned along the $Y$ axis, symmetrically aligned around the origin and directed towards one another. The first camera is tasked with reconstructing the original image, while the second models the mirror reflection. The photograph is thus conceptualized as a translucent tracing paper sheet, embedded within the 3D spatial context. The reflection can be effectively represented by horizontally flipping the image, denoted as $\mathcal{M}(I)$. This mirror-camera setup enhances the fidelity of the generated reflections, providing a robust solution for accurately capturing visual elements.
The MiraGe is initialized according to equation Eqn. 8 and utilizes a cost function: $\mathcal{L}(I)+L(\mathcal{M}(I))$. We simultaneously model both the image and its mirrored reflection, as shown in the second row in Fig. 6.
Figure 13: MiraGe allows us to produce realistic modifications of small details like changing human facial expressions.
After thorough experimentation, we find that our model, Amorphous-MiraGe, utilizing a mirror camera, achieves state-of-the-art reconstruction results. This model demonstrates significant advantages over alternative methods in terms of both performance and outcome quality.
Editability
The ability to manipulate Gaussians based on their spatial positioning empowers MiraGe to effectively edit 2D images. When utilizing a mirror camera, the quality of the resulting images is sufficiently high, enabling the parameterization and animation of Gaussians to significantly reduce artifacts. Our findings demonstrate that our model facilitates the animation of both segmented objects and entire scenes. Users can create manual animation, or leverage automated processes using physics engines like Taichi_elements or Blender (8, 9). To incorporate MiraGe with the 2D physics engine, we use 2D-MiraGe (Fig 8).
We argue that the Graphite-inspired model allows the creation of attractive compositions made of multiple images that effectively present the positive attributes of the layered structure, like Graphite, through the strategic positioning of Gaussians.
4 Experiments
We split the experimental section of our paper into two main parts. First, we demonstrate that our approach achieves high-quality 2D reconstruction by comparing it with existing models. Second, we highlight the versatility of our method in image editing both full scenes (Fig. 10) and selected objects (Fig. 1, 7), presenting examples of user-driven modifications and demonstrations involving physical simulations (Fig. 8, 9).
Figure 14: MiraGe model allows modifications in 3D space, but the model is limited by 2D images, which was used in training. When we move some elements from the foreground, we are not able to see the background since the model only reconstructs objects. After modification, we can see artifacts on the hind paw of the presented animals.
Reconstruction quality
To assess image quality, we utilized two established metrics, i.e., PSNR and MS-SSIM (Wang et al., 2003). These metrics measure the distortion between the reconstructed images and the ground truth.
Our image reconstruction assessment utilizes two widely-recognized datasets. Specifically, we employ the Kodak dataset [^2] https://r0k.us/graphics/kodak/, which includes 24 images at a resolution of $768\times 512$, alongside the DIV2K validation set (Agustsson & Timofte, 2017), which involves 2×2$\times$2 × bicubic downscaling and comprises 100 images with sizes ranging from $408\times 1020$ to $1020\times 1020$. The dataset was selected to facilitate direct comparison with the work of GaussianImage.
As a baselines we use competitive INR methods GaussianImage (Zhang et al., 2024), SIREN (Sitzmann et al., 2020b), WIRE (Saragadam et al., 2023b), I-NGP (Müller et al., 2022), and NeuRBF (Chen et al., 2023).
In Tab. 1, we demonstrate the performance outcomes of different methods on the Kodak and DIV2K datasets. We see that our proposition outperforms the previous solutions on both datasets. The quality measured by both metrics shows significant improvement compared to all the previous approaches. Fig. 11 illustrates a general trend observed during training in contest of image reconstruction. The selection of hyperparameters, including the number of iterations, was inspired by the principles of 3DGS. We provide ablation studies and extensive numerical analyses in the appendix for further insights.
It is important to note that although our model takes longer to train, it quickly achieves better results than GaussianImage. The trend we observe is illustrated in Fig. 11, which also includes the number of initial Gaussians, indicating how densely the space has been filled. We see a clear upward trend in performance as the density of the Gaussian initialization increases.
Due to our particular focus on animation, we analyzed FPS trends to benchmark real-time performance. Fig. 12 shows that while our model introduces a higher number of parameters, leading to a decrease in FPS compared to GaussianImage, it maintains the ability to render animations in real time
Manual modification
MiraGe allows for manual manipulation of 2D images. By leveraging GaMeS parameterization, each Gaussian component is represented as a triangle. Vertices can then be independently adjusted and moved within 3D space, enabling flexible image modification (Fig. 27).
We demonstrate examples of modifications using datasets such as DIV2K, Kodak, and Animals [^3] https://www.kaggle.com/datasets/alessiocorrado99/animals10. Additionally, we generated our own 2D images using DALL-E 3 to illustrate the benefits of our method. We can obtain modifications of small details like changing fingers position (Fig. 1), human facial expressions (Fig. 13) or dog poses (Fig. 14). As MiraGe can trained in a 3D context, we can implement modifications in the third dimension to create the illusion of 3D transformation (Fig. 36).
It is crucial to note that when we displace elements from the foreground, the background remains unseen because the model only reconstructs the objects. This is demonstrated in Fig. 14, where artifacts are apparent on the hind paw of the depicted animal.
Table 1: Quantitative comparison with various baselines in PSNR and MS-SSIM. MiraGe gives state of the art results. Model-$x$ denotes that the model was initialized with $x$ Gaussians.
Kodak dataset DIV2K dataset
PSNR $\uparrow$ MS-SSIM $\uparrow$ PSNR $\uparrow$ MS-SSIM $\uparrow$
WIRE (Saragadam et al., 2023a) 41.47 0.9939 35.64 0.9511
SIREN (Sitzmann et al., 2020a) 40.83 0.9960 39.08 0.9958
I-NGP (Müller et al., 2022) 43.88 0.9976 37.06 0.9950
NeuRBF (Chen et al., 2023) 43.78 0.9964 38.60 0.9913
3DGS (Kerbl et al., 2023) 43.69 0.9991 39.36 0.9979
GaussianImage-70k (Zhang et al., 2024) 44.08 0.9985 39.53 0.9975
GaussianImage-100k* (Zhang et al., 2024) 38.93 0.9948 41.48 0.9981
MiraGe-70k (our) 57.41 0.9998 53.22 0.9996
MiraGe-100k (our) 59.52 0.9999 54.54 0.9998
Physics application in MiraGe
Using 2D-MiraGe we can express Gaussian components with a 2D point cloud. Therefore, we can use MPM (Hu et al., 2018) based physics engine implemented, for example, in Taichi__\__elements. This high-performance physics engine supports multiple materials, including elastic objects and sand. We use inspiration from GASP (Borycki et al., 2024) and train simulation on 2D points then use physical deformation on triangle soup. In Fig. 8, we present simulation results obtained using Taichi_elements. As we can see, we can add physical properties to 2D objects. On the other hand, using Amorphous-MiraGe or Graphite-MiraGe we can use Blender and modify directly parameterized flat 3D Gaussian (Fig. 9).
5 Conclusion
In this paper, we introduce MiraGe that uses flat 3D Gaussian components to model 2D images. MiraGe gives state-of-the-art reconstruction quality and simultaneously allows image manipulation. Furthermore, we can modify photos on a plane (Fig. 8) and in 3D space (Fig. 10, 9). In consequence, we obtain the illusion of 3D-based modifications. Furthermore, we can combine our solution with a physics engine to obtain realistic motion in the image. Conducted experiments show that MiraGe is applicable in many different scenarios and produces high-quality simulations.
Limitation
It is crucial to note that the model is not generative, so improper adjustment of Gaussian positions can cause gaps in the image (e.g. a missing dog’s paw 14). While the model can produce realistic changes, significant modification may introduce visual artifact. Moreover, our model requires encoding more parameters than GaussianImage to achieve high-quality image reconstruction for animation. Addressing this trade-off will be a focus of our future work.
Social impact
The model can be applied to generate novel image transformations used to dataset augmentation, facial recognition (Sanil et al., 2023), distortion correction, and visualizing architectural designs (Gerstweiler et al., 2018), as shown in Fig. 10. In the field of medical imaging, it is utilized for refining anatomical models and enhancing the accuracy of surgical simulations and diagnostic tools (Lin et al., 2023). Additionally, in 2D computer games, the model facilitates more realistic animations by incorporating physics-based effects (Mohd et al., 2023).